Determine 2 {F}. 5s? +17s +18 F(s) = (s + 2)2 (8+6) Click here to view the table of Laplace transforms. ul. Ella qº{F} =
Answer #1

+ + 582 +17s +18 Given function F(s) = (s+2) (s+6) 582 +17s +18 B (s+2) (s+6) (s+2) (s+2) (s+6) 55* +175+18= Aſs+2)(s+6)+B(